THE NEAR INFRARED Y^{2}Σ^{+} − X^{2}Π TRANSITION OF COPPER SELENIDE

\underline{LEAH O’BRIEN} and AMANDA LAMBETH, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The near infrared electronic transition of CuSe has been observed for the first time. The spectrum of the Y^{2}Σ^{+} − X^{2}Π transition, labeled by analogy with the CuO and CuS near infrared electronic transitions, was recorded with the Fourier transform spectrometer associated with the McMath-Pierce Solar Telescope, Kitt Peak, AZ. The CuSe molecules were produced in a King-type carbon tube furnace operating at 1800 C and at a total pressure of approximately 400 torr. The electronic term energy and vibrational constants for both states will be presented.